A rock falls from a tower that is 73.5 m high. As it is falling, its height is given by the formula h=73.5 -4.9t^2, with t in seconds. How many seconds will it take for the rock to hit the ground? (h=0)?
Answer by Alan3354(30993)

h=73.5 -4.9t^2 = 0
t^2 = 73.5/4.9
t^2 = 15
t =~ 3.873 seconds